zarkam21:

A boat floats south on the Amazon River at a speed of 6 m/s. The boat and passengers have a combined mass of 540 kg. a. What is the momentum of the boat? (Remember, momentum is a vector quantity.) b. The passengers paddle backward, reducing the speed to 4 m/s. What was the impulse applied to the boat by the paddling?

Vocaloid:

a) p = mv as usual

zarkam21:

P=540*6 p=3240 kg*m/s

Vocaloid:

good, since momentum is a vector we also need to specify the direction so p=3240 kg*m/s south

Vocaloid:

for b) impulse = change in momentum = ?

zarkam21:

2?

zarkam21:

would it be 6-4

Vocaloid:

that's the change in velocity what about the change in momentum?

zarkam21:

p=540*4 p=2160 3240-2160=1080

zarkam21:

1080 was the change in momentum

Vocaloid:

good, and since the impulse is in the opposite direction as the initial momentum we say the impulse was 1080 kg m/s north

zarkam21:

At a race track, a car of mass 1150 kg crashes into a concrete wall at a speed of 85 m/s. a. If the car comes to a stop when it hits the wall, what is the magnitude of the impulse applied to the car? b. The car crashes into the wall, stopping in 0.1 seconds. What force is applied to the car?

Vocaloid:

a) impulse is change in momentum (magnitude means we can disregard direction for now)

zarkam21:

p=1150*85 p=97750

Vocaloid:

good (units are kg m/s as usual) for b) impulse = f * delta t, calculate f

zarkam21:

97750=f*0.1

Vocaloid:

good, solve for f

zarkam21:

977500N

Vocaloid:

good

zarkam21:

Bill forgets to put his car in park, and it starts rolling forward. When it is moving at a speed of 2 m/s, it collides with Tanya's car, which is stationary. Bill's car bounces backward at a speed of 1.2 m/s. Bill's car has a mass of 1100 kg, and Tanya's car has a mass of 800 kg. a. What is the total momentum of the system before the collision? b. What is the total momentum of the system after the collision? c. What is the velocity of Tanya's car after the collision?

Vocaloid:

a) total momentum = initial momentum of Billy's car + initial momentum of Tanya's car = ?

zarkam21:

i combine the masses first right?

zarkam21:

oh okay so do it separately

Vocaloid:

hint: initially, only bob's car is moving so it's just bob's car mass * velocity

zarkam21:

Bob p=mv p=1100*2 p=2200 kg m/s

zarkam21:

Tanya p=mv p=800*2 p=1600 kg m/s

Vocaloid:

good, and the direction is "forward" so 2200 kg m/s forward b) momentum is conserved so final momentum = initial momentum for part c) set bob's momentum equal to tanya's and solve for tanya's velocity

zarkam21:

2200-1600=600

Vocaloid:

tanya's car is not moving at the beginning so you do not need to add tanya's momentum for parts a and b.

zarkam21:

okay cool so just bob's momentum

zarkam21:

okay now for b

zarkam21:

i combine the two masses and then calculate momentum right

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid good, and the direction is "forward" so 2200 kg m/s forward b) momentum is conserved so final momentum = initial momentum for part c) set bob's momentum equal to tanya's and solve for tanya's velocity \(\color{#0cbb34}{\text{End of Quote}}\) the answers for part a and part b should be the same because momentum is conserved.

zarkam21:

so it would just be 2200 for B

Vocaloid:

yes

zarkam21:

so b.) 2200 kg m/s forward c.)2200=960

Vocaloid:

for part c) the statement you have given is not logical, 2200 is not equal to 960 set Bob's final velocity * Bob's car mass = Tanya's final velocity * Tanya's car mass and solve for tanya's velocity

zarkam21:

1100=1.2*800

zarkam21:

wait

zarkam21:

1100=v(800)

zarkam21:

1.4?

Vocaloid:

bob's final momentum = 1100 * 1.2 = tanya's final momentum = 800 * v solve for v

zarkam21:

1320=800*v v=1.65

Vocaloid:

good, 1.65 m/s forward

zarkam21:

thank you so much!